# Stability analysis of very high order minimization-based and Taylor-based embedded boundary treatments of discontinuous Galerkin for hyperbolic equations

### v1

@@ Abstract
In this paper, we present a stability analysis of very high order embedded boundary methods, specifically the Reconstruction for Off-site Data (ROD) and Shifted Boundary (SB) methods, coupled with a discontinuous Galerkin discretization for the linear advection equation. In unfitted configurations, these methods impose consistent modified boundary conditions on the computational boundary. Due to the high algebraic complexity of very high order schemes, the stability is studied by visualizing the eigenspectrum of the discretized operators. A recent study on the SB method demonstrated that its Taylor expansion can be formulated as a direct polynomial correction. In this work, we prove that the ROD minimization problem admits an analogous polynomial correction. This unified perspective provides significant benefits: algorithmically, it greatly simplifies the implementation of ROD by eliminating the need for linear system inversions at each iteration; mathematically, it enables a rigorous stability study. For completeness, a side-by-side stability analysis of the ROD and SB methods is presented for polynomials up to degree 6. Furthermore, due to the stability restrictions of embedded methods for hyperbolic problems, a coupling with both explicit and implicit time integration is investigated. A set of numerical experiments confirms the findings of the stability study.
Introduction
This work focuses on the stability analysis of very high order discontinuous Galerkin (DG) methods [reed1973triangular, cockburn2001runge] with embedded boundary treatments, for the linear advection equation in one space dimension:
@@ Introduction
where $u$ is the conservative advected variable, while the advection speed is equal to one. $x$ and $t$ represent the space and time variables, respectively.
High order methods have the potential in achieving very accurate solutions with reduced computational cost with respect to low order methods [wang2013high]. However, ensuring consistency and stability for high order boundary conditions remains a very challenging task, especially when dealing with complex curved geometries. When dealing with non-trivial boundaries, accurately representing geometric features of the computational domain is as important as the numerical method itself [bassi1997high]. As a matter of fact, the global accuracy of the numerical solution is given by both the accuracy of the numerical method and the precision used for the geometry representation.
A widespread approach in high order numerical simulations is the isoparametric approach, whose goal is to use high order polynomial reconstructions to approximate the given geometry, with the same precision of the numerical method used to approximate the problem [bassi1997high]. This method achieves optimal convergence rates and improved accuracy in the presence of curved boundaries, and allows one to discretize complex geometries with very coarse meshes. However, the generation of high order curvilinear meshes remains a real challenge: nonlinear mappings, complex quadrature rules, and a higher number of required degrees of freedom are just a few of the tasks that one needs to address. Furthermore, additional difficulties related to mesh generation [persson2009curved] arise, which needs advanced techniques to ensure mesh validity and quality. Despite recent progress, dealing with curvilinear meshes remains a multifaceted task for realistic applications [coppeans2025aerodynamic].
The main goal of embedded boundary methods is to reduce the effort related to mesh generation and mesh motion by only employing simple, fixed, easily generated meshes, over which embedded boundaries can move freely. However, contrary to body-fitted meshes, this treatment introduces a geometrical error proportional to the mesh size $h$. In this case, the numerical challenge is no longer related to mesh generation but to the development of accurate and stable embedded boundary conditions. The main idea behind these methods was introduced in the pioneering work of Peskin [peskin2002immersed]. In the last fifty years, there has been a huge effort to develop unfitted methods with the objective of minimizing computational costs related to mesh generation and motion. Some of these works are found in [mittal2005immersed, fadlun2000combined, hansbo2002unfitted, fedkiw1999non, berger2012progress, monasse2012conservative, coco2013finite], and references therein. Although most of these methods are first and second order accurate, there has been a recent interest in developing high order embedded boundary methods, to be coupled with arbitrary high order methods used to discretize the physical problem. Recent works in this direction can be found in [tan2010inverse, atallah2022high, giuliani2022two].
This paper contributes to the analysis and development of arbitrary high order embedded boundary treatments applied to hyperbolic partial differential equations (PDEs). In this work, we propose to study numerically the stability analysis of very high order DG methods coupled with the Reconstruction for off-site Data (ROD) method [ciallella2024very, santos2024very], a minimization-based boundary treatment, and compare it with the analysis of the Shifted Boundary (SB) polynomial correction [ciallella2025stability], a Taylor-based boundary treatment, for problems with fully unfitted boundary configurations: when the distance between the exact boundary and the computational one is of the order $h$. The shifted boundary polynomial correction was introduced in [ciallella2023shifted] to simulate hyperbolic PDEs, as an efficient alternative to the high order shifted boundary method [atallah2022high], introduced previously for elliptic PDEs. In particular, in [ciallella2023shifted], it was shown that it is possible to rewrite the truncated Taylor expansion of the SB method as a polynomial correction, given by the evaluation of the internal polynomial on the real boundary point. This simplifies the development, which no longer needs the computation of high order derivatives terms, and also made easier the numerical stability analysis performed in [ciallella2025stability]. Some applications of the shifted boundary method can be found in [main2018shifted, atallah2022high, visbech2025spectral, ciallella2020extrapolated, ciallella2022extrapolated, BOSCHERI2025114215].
The ROD method was developed, in finite volume [costa2018very, costa2019very] and in DG [santos2024very, ciallella2024very], to impose high order boundary conditions when curved boundaries are discretized by simplicial meshes (triangles or tetrahedrons), i.e. when the distance between the exact boundary and the computational one is of the order $h^{2}$. The ROD method, for DG methods, consists in imposing consistent embedded boundary conditions by building a modified polynomial that is as close as possible to the internal one, and satisfies exactly the given boundary condition on the real boundary. Therefore, the coefficients of the ROD polynomial are simply defined as the outcome of a constrained minimization problem. The goal of this work is to study the eigenspectrum of the discretized operators of high order DG approximations of linear hyperbolic equations, when introducing the ROD embedded boundary treatment. Studying the eigenspectrum of the discretized operators to analyze complex boundary treatments has shown to be very effective [M2AN_2015__49_1_39_0, yang2025inverse, ciallella2025stability], especially for high order methods, for which analytical studies become very cumbersome. However, the definition of the ROD polynomial as the result of a constrained minimization problem does not make straightforward its integration in a stability analysis. To do so, it would be necessary to write the ROD boundary treatment in matrix form, as it was done in [ciallella2025stability] for the SB method.
Therefore, the first contribution of this work is to show that the ROD minimization boundary treatment can be explicitly recast as a polynomial correction, similarly to the SB polynomial correction [ciallella2023shifted] with an additional multiplicative factor. Therefore, the ROD method boils down to a polynomial correction, which explicitly depends on a set of modified basis functions that can be computed once and stored for the whole simulation time. On one side, this formulation clearly simplifies the development of the ROD boundary treatment, making it more efficient to implement when compared to solving explicitly the minimization problem at each time step. On the other, recasting the ROD approach as a polynomial correction allows us to write it straightforwardly in matrix form, which is fundamental to rigorously study its stability properties when analyzing the eigenspectrum of discretized operators. The analysis of the ROD minimization-based method is presented for very high order methods and, to give a larger overview, it is compared with the results obtained with the SB Taylor-based approach [ciallella2025stability]. Recent works on the ROD method [ciallella2024very, santos2024very] have set up the minimization problem using the squared Euclidean norm; herein, we also show that, by using the squared $L^{2}$ norm, it is possible to achieve a scheme that has improved stability properties. In addition to that, we prove that all these versions of the ROD method can be recast as polynomial corrections, which differ only in the definition of the multiplicative factor. The results of the analysis show that, for polynomials of degree $p\leq 4$, it is possible to derive a ROD method that is stable for explicit time integration, under a standard CFL condition. This is an important improvement when compared to the SB method, which shows a restriction on the CFL also for lower order polynomials. In the second part of the analysis, it is investigated whether implicit time integration can improve the stability properties of the embedded methods considered herein. Although the stability regions are enlarged for implicit time integration, we observe that for very high order methods all considered approaches present lower bound conditions on the CFL, which needs to be high enough to ensure unconditional stability. All results obtained through the proposed analysis are validated through a large set of numerical experiments where several boundary configurations are considered, and with accuracy varying between second and seventh order.
@@
The rest of the paper is organized as follows. In section 1 the high order DG discretization for the linear advection equation is presented. In section 2 the notation of the embedded boundary methods is introduced. Section 3 describes the high order Taylor-based embedded boundary treatment, namely the SB polynomial correction. In section 4, we show that several variations of the ROD minimization-based method can be recast as a polynomial correction. The stability analysis for all embedded boundary methods are presented in section 5. The analysis is applied first to DG with standard periodic boundary conditions in section 5.1 to show that standard CFL conditions are retrieved. Then, the DG methods with the considered embedded boundary method and explicit time integration are studied in section 5.2 and section 5.3; while their coupling with implicit time integration are considered in section 5.4 and section 5.5. Numerical tests that validate all the theoretical results are presented in section 6, and some conclusions are drawn in section 7.
1 High order discontinuous Galerkin for the linear advection equation
For simplicity, this section describes the considered high order discontinuous Galerkin method in one space dimension. Let us consider a one-dimensional computational domain $\Omega$ discretized into $N_{e}$ non-overlapping elements $\Omega_{i}$, such that $\Omega=\bigcup_{i=1}^{N_{e}}\Omega_{i}$. Without lack of generality, the cell $\Omega_{i}=[x_{i-1/2},x_{i+1/2}]$ has a fixed size of $\Delta x$. The numerical solution $u$ is approximated by $u_{h}$, which belongs to the global discrete space $V_{h}$ defined as:
$V_{h}:=\bigoplus_{i=1}^{N_{e}}V_{i},\quad\text{where}\quad V_{i}:=\mathbb{P}^{p}(\Omega_{i}).$ (2)
@@ 1 High order discontinuous Galerkin for the linear advection equation
$V_{i}$ is the local approximation space on element $\Omega_{i}$, given by the space of polynomials of degree $p\geq 0$. Hence, the numerical solution is a piece-wise polynomial that is discontinuous across element interfaces.
@@
Inside each element $\Omega_{i}$, the approximation $u_{h}$ is expressed as a linear combination of basis functions $\{\varphi_{n}\}_{n=0}^{p}$ of the space $V_{i}$:
@@
where $u_{i,n}(t)$ are the degrees of freedom of cell $\Omega_{i}$.
For completeness, the linear advection equation with a source term will be considered for the space discretization:
@@
Below, the notation $(x,t)$ will be dropped for compactness.
The elemental semi-discrete DG weak formulation is written by projecting (4) and integrating by parts,
$\int_{\Omega_{i}}\varphi^{i}_{m}\frac{\text{d}u_{h}}{\text{d}t}\,\text{d}x-\int_{\Omega_{i}}\partial_{x}\varphi^{i}_{m}u_{h}\,\text{d}x+\int_{\partial\Omega_{i}}\varphi^{i}_{m}\hat{u}(u_{h}^{-},u_{h}^{+})n\,\text{d}\Sigma=\int_{\Omega_{i}}\varphi^{i}_{m}s(x)\,\text{d}x,\quad\forall\varphi^{i}_{m}\in V_{i}.$ (5)
$n$ is the outward normal, and $\hat{u}(u_{h}^{-},u_{h}^{+})$ is a numerical flux that depends on the left $(\cdot)^{-}$ and right $(\cdot)^{+}$ reconstructed interface values. In one-dimension, equation (5) can be simply rewritten as
@@
The numerical flux at the element interfaces is computed using a simple upwind flux $\hat{u}(u_{h}^{-},u_{h}^{+})=u_{h}^{-}$, which is equal to the local Lax-Friedrics numerical flux with positive advection speed.
When replacing (3) into equation (6), the semi-discrete problem for a generic element $\Omega_{i}$ can be written in matrix form as follows
@@
where $\mathbf{u}_{i}=[u_{i,0},\ldots,u_{i,p}]^{T}$ is the vector of degrees of freedom belonging to element $\Omega_{i}$. The mass and stiffness matrices are defined as,
$M_{i}=\int_{\Omega_{i}}\varphi^{i}_{m}\varphi^{i}_{n}\,\text{d}x,\qquad K^{s}_{i}=\int_{\Omega_{i}}\partial_{x}\varphi^{i}_{m}\varphi^{i}_{n}\,\text{d}x,\quad\forall\varphi^{i}_{m},\varphi^{i}_{n}\in V_{i},$ (8)
while interface matrices are defined as,
$K^{R}_{i}=\varphi_{m}^{i}(x_{i+1/2})\varphi_{n}^{i}(x_{i+1/2}),\qquad K^{L}_{i}=\varphi_{m}^{i}(x_{i-1/2})\varphi_{n}^{i-1}(x_{i+1/2})$ (9)
where $\varphi_{m}^{i}(x_{i+1/2})$ represents the $m$-th basis of the element $\Omega_{i}$ evaluated in $x_{i+1/2}$.
To trace the stability region of the methods from the eigenspectrum, it is necessary to write the system in a global matrix form of the kind
@@
where $\mathbf{M}$ and $\mathbf{K}$ represent the global mass and stiffness matrices, respectively. $\mathbf{U}=[\mathbf{u}_{1},\ldots,\mathbf{u}_{N_{e}}]^{T}$ is the global vector of degrees of freedom.
To derive the eigenvalue problem for the explicit time integration of the semi-discrete system (10), we consider the amplification factor of arbitrary high order Deferred Correction (DeC) time integration schemes [dutt2000spectral, micalizzi2024new] of order $p+1$, which gives
@@
where $\mathbf{U}^{n}=\mathbf{U}(t^{n})$ and $\mu=\lambda\Delta t$, with $\Delta t$ the time step and $\lambda$ the eigenvalue of semi-discrete operator $\mathbf{M}^{-1}\mathbf{K}$. In this work the polynomial degree for the space discretization is taken up to $p=6$ and it is coupled with DeC up to order 7. Therefore, the stability region of the considered fully discrete problem is $|z(\mu)|\leq 1$, where $z(\mu)=\sum_{k=0}^{p+1}\frac{\mu^{k}}{k!}$.
To perform the analysis is therefore needed to recast the embedded boundary treatment in the general matrix form $K\mathbf{u}$. Although this is trivial for the SB correction as it was previously shown in [ciallella2025stability] and recalled in section 3, this is not straightforward for the minimization-based ROD techniques which need careful manipulation to be written in such a form. In section 4, the ROD minimization problem is explicitly solved to show that this class of methods as well can be written as a special polynomial correction, as it was done for the SB method. This allows one to make a link between the two families of methods and, more interestingly allows one to perform a rigorous stability analysis.
@@ 2 Notation of embedded boundary methods
Figure 1: Mesh configuration: tessellation of one-dimensional elements with a boundary condition (red circle) imposed at a distance $d$ from the left-most interface of the domain, which is considered as the computational boundary of the problem.
The aim of this work is to explore the impact of the ROD minimization-based treatment and, to have a larger overview, to compare it to the SB Taylor-based treatment when the distance $d$ from the real boundary to the computational one is of the order of the mesh size $\Delta x$. This is a classical configuration that can occur when dealing with numerical simulations with boundaries treated as immersed starting from a Cartesian mesh (see figure 1 for more details). In multi-dimensional configurations the mapping between the real boundary and the computational boundary is not unique and can be prescribed using different approaches: closest-point, level-set, projections. However, in one space dimension, the mapping between the two is uniquely identified, which simplifies the analysis.
Consider $\bar{\mathbf{x}}$ the position of a point on the real boundary, while $\tilde{\mathbf{x}}$ the position of the corresponding computational boundary point. The mapping between the two is given by $\bar{\mathbf{x}}=\tilde{\mathbf{x}}+\mathbf{d}$, where $\mathbf{d}$ is the distance vector. Considering $\mathbf{n}_{e}$ the unit outward normal of the boundary face, the signed distance can be defined as the following scalar product:
@@
Following this notation, the distance is allowed to vary within the range $d\in[-\Delta x,\Delta x]$: meaning, the real boundary can be either outside the first mesh element (when $d<0$), or inside the first mesh element (when $d>0$), as shown in figure 2. In a one-dimensional case, the mapping can be rewritten in a simple scalar form $\bar{x}=\tilde{x}+d$.
As a standard interface, the boundary condition is imposed at the computational boundary through a numerical flux. For an upwind flux, the left boundary flux of the domain is given by $\hat{u}(v_{h},u^{+}_{h})=v_{h}$, where $v_{h}$ corresponds to the considered high order embedded boundary reconstruction.
The semi-discrete evolution equation (6) with an upwind numerical flux for the boundary element $\Omega_{b}=[x_{1/2},x_{3/2}]$ simply reads
$\begin{split}&\int_{\Omega_{b}}\varphi^{i}_{m}\frac{\text{d}u_{h}}{\text{d}t}\,\text{d}x-\int_{\Omega_{b}}\partial_{x}\varphi^{i}_{m}u_{h}\,\text{d}x+\\ &\qquad\qquad\bigl[\varphi^{i}_{m}(x_{3/2})u_{h}(x_{3/2})|_{i}-\varphi^{i}_{m}(x_{1/2})v_{h}(x_{1/2})\bigr]=\int_{\Omega_{b}}\varphi^{i}_{m}s(x)\,\text{d}x.\end{split}$ (13)
[…]
Figure 2: Boundary position: the real boundary position can be either outside the first mesh element (above), or inside the first mesh element (below). In the first case, the distance will be taken as negative $d<0$, while for the second one the distance is taken as positive $d>0$.

### v2

# Minimization-based embedded boundary methods as polynomial corrections: a stability study of discontinuous Galerkin for hyperbolic equations

@@ Abstract
This work establishes a novel, unified theoretical framework for a class of high order embedded boundary methods, revealing that the Reconstruction for Off-site Data (ROD) treatment shares a fundamental structure with the recently developed shifted boundary polynomial correction [Ciallella, M., et al. (2023)]. By proving that the ROD minimization problem admits an equivalent direct polynomial correction formulation, we unlock two major advances. First, we derive a significant algorithmic simplification, replacing the solution of the minimization problem with a straightforward polynomial evaluation, thereby enhancing computational efficiency. Second, and most critically, this reformulation enables the first stability result for the ROD method when applied to the linear advection equation with discontinuous Galerkin discretization. Our analysis, supported by a comprehensive eigenspectrum study for polynomial degrees up to six, characterizes the stability region of the new ROD formulation. The theoretical findings, which demonstrate the stability constraints, are validated through targeted numerical experiments.
1 Introduction
high order discontinuous Galerkin (DG) methods [28, 10] offer a powerful framework for achieving accurate solutions with improved computational efficiency [34]. However, their application to problems with complex geometries presents a significant challenge, as the accuracy of the final numerical solution is intrinsically linked to the precision of the geometry representation [2]. The isoparametric approach, which uses high order polynomial mappings to represent curved boundaries, can provide optimal convergence [2]. However, the practical difficulties associated with generating and validating high order curvilinear meshes remain relevant hurdles for realistic applications [25, 12]: nonlinear mappings, complex quadrature formulas, and mesh quality control are the major issues to be addressed.
Embedded (or immersed) boundary methods present an alternative paradigm, aiming to circumvent these mesh generation challenges by using simple, fixed background grids over which complex boundaries can move freely. This shift transfers the primary difficulty from mesh generation to the formulation of accurate and stable boundary conditions, as the method introduces a geometric error proportional to the mesh size $h$. Following the pioneering immersed boundary work of Peskin [26], extensive development has produced many methods for fluid and solid mechanics [16, 20, 17, 23, 3, 24, 11, 27]. While historically many were low-order, recent years have seen growing interest in high order embedded treatments compatible with high order solvers [31, 1, 18, 8, 9, 4, 33], which better adapt to modern computer architectures.
This work advances this field by addressing a specific gap: the stability of minimization-based high order embedded methods, whose analysis has been hindered by their implicit formulation. We focus on the Reconstruction for Off-site Data (ROD) method [13], initially developed to impose high order conditions for curved boundaries on simplicial meshes within finite volume [13, 14] and DG frameworks [29, 7]. In a DG context, ROD operates by constructing a modified boundary polynomial via a constrained minimization problem, which is as close as possible to the internal polynomial and exactly satisfies the true boundary condition. This modified polynomial is then used to impose consistent boundary conditions on a computational boundary, which does not match the real boundary. The central analytical obstacle is that this minimization-based definition cannot be directly employed for stability investigation. Therefore, our first and fundamental contribution is to prove that the ROD formulation is mathematically equivalent to an explicit polynomial correction, akin in structure to the shifted boundary polynomial correction [8, 6]. This equivalence reveals ROD to be part of a broader unified framework for embedded treatments. Practically, it yields a significant algorithmic simplification, replacing the solution of a minimization problem with the application of a precomputable correction. This reformulation enables our second contribution: a rigorous stability analysis of high order DG methods coupled with the ROD treatment for the one-dimensional linear advection equation
@@ 1 Introduction
where $u$ is the conservative variable and the advection speed is equal to one. By recasting the method in polynomial correction form, we can straightforwardly construct the discretized operator and study its eigenspectrum, an established technique for analyzing complex high order boundary treatments [32, 35, 6], which is made even easier by the local structure of DG. We extend the analysis beyond the standard ROD formulation, based on the Euclidean norm, to consider an $L^{2}$-norm variant, demonstrating how the choice of norm influences stability. Our findings reveal that, for polynomial degrees $p\leq 4$, a stable explicit scheme can be constructed under a standard CFL condition, an improvement over related embedded methods [6]. For very high orders, we characterize the stability limitations and show that the implicit time integration can improve them, although with specific CFL constraints: in particular, we show that the CFL has to be high enough to ensure stability, which does not make it optimal if one studies unsteady problems. All theoretical results are validated through comprehensive numerical experiments spanning multiple boundary configurations and orders of accuracy from second to seventh.
@@
The rest of the paper is organized as follows. In section 2 the high order DG discretization for the linear advection equation is presented. In section 3 the notation of the embedded boundary methods is introduced. Section 4 describes the high order SB polynomial correction, which is used to make a link with the novel ROD polynomial corrections. In section 5, we show that several variations of the ROD minimization-based method can be recast as a polynomial correction, similar to the one described in section 4. The stability analysis for all embedded boundary methods are presented in section 6. The analysis is applied first to DG with standard periodic boundary conditions in section 6.1 to show that standard CFL conditions are retrieved. Then, the DG methods with the considered embedded boundary method and explicit time integration are studied in section 6.2; while their coupling with implicit time integration are considered in section 6.3. Numerical tests that validate all the theoretical results are presented in section 7, and some conclusions are drawn in section 8.
2 Discontinuous Galerkin for linear advection
This section introduces the one-dimensional DG formulation that serves as the foundation for the analysis of the embedded boundary treatments considered in this work. Consider a one-dimensional computational domain $\Omega$ partitioned into $N_{e}$ non-overlapping cells $\Omega_{i}$, such that $\Omega=\bigcup_{i=1}^{N_{e}}\Omega_{i}$. For simplicity, each cell $\Omega_{i}=[x_{i-1/2},x_{i+1/2}]$ has a fixed size of $\Delta x$. The discrete solution $u_{h}$ resides in the broken polynomial space
$V_{h}:=\bigoplus_{i=1}^{N_{e}}V_{i},\qquad\quad V_{i}:=\mathbb{P}^{p}(\Omega_{i}),$ (2)
@@ 2 Discontinuous Galerkin for linear advection
where $V_{i}$ denotes the local space of polynomials of degree $p\geq 0$ on cell $\Omega_{i}$. Thus, $u_{h}$ is a piece-wise polynomial that may be discontinuous at cell interfaces.
@@
Within each element $\Omega_{i}$, the numerical solution $u_{h}$ is expressed in terms of local basis functions $\{\varphi_{n}\}_{n=0}^{p}$ of the space $V_{i}$:
@@
where $u_{i,n}(t)$ are the time-dependent modal coefficients of cell $\Omega_{i}$.
To be aligned with the numerical experiments, the model problem under consideration is the linear advection equation with a source term:
@@
which will be used to derive the semi-discrete form. For brevity, the explicit dependence on $x$ and $t$ is often omitted below.
The DG weak formulation in the semi-discrete form for cell $\Omega_{i}$ is obtained by projecting (4) and applying integration by parts,
$\int_{\Omega_{i}}\varphi^{i}_{m}\frac{\text{d}u_{h}}{\text{d}t}\,\text{d}x-\int_{\Omega_{i}}\partial_{x}\varphi^{i}_{m}u_{h}\,\text{d}x+\int_{\partial\Omega_{i}}\varphi^{i}_{m}\hat{u}(u_{h}^{-},u_{h}^{+})n\,\text{d}\Sigma=\int_{\Omega_{i}}\varphi^{i}_{m}s(x)\,\text{d}x,\quad\forall\varphi^{i}_{m}\in V_{i},$ (5)
[…]
@@
Substituting the expansion (3) into (6) yields the semi-discrete system for element $\Omega_{i}$ in matrix-vector form:
@@
where $\mathbf{u}_{i}=[u_{i,0},\ldots,u_{i,p}]^{T}$ is the vector of local degrees of freedom. The elemental mass matrix $M_{i}$, the stiffness matrix $K^{s}_{i}$, and the interface flux matrices are defined as,
$M_{i}=\int_{\Omega_{i}}\varphi^{i}_{m}\varphi^{i}_{n}\,\text{d}x,$ $\qquad K^{s}_{i}=\int_{\Omega_{i}}\partial_{x}\varphi^{i}_{m}\varphi^{i}_{n}\,\text{d}x,$ (8)
$K^{R}_{i}=\varphi_{m}^{i}(x_{i+1/2})\varphi_{n}^{i}(x_{i+1/2}),$ $\qquad K^{L}_{i}=\varphi_{m}^{i}(x_{i-1/2})\varphi_{n}^{i-1}(x_{i+1/2}),$ (9)
where $\varphi_{m}^{i}(x_{i+1/2})$ represents the $m$-th basis of cell $\Omega_{i}$ in $x_{i+1/2}$.
The stability analysis is performed by examining the eigenspectrum of the semi-discrete system. To proceed, we first assemble the local DG formulation into a global system of ordinary differential equations:
@@
where $\mathbf{M}$ is the block-diagonal global mass matrix, $\mathbf{K}$ is the global stiffness matrix incorporating both interior stiffness and boundary conditions, and $\mathbf{U}=[\mathbf{u}_{1},\ldots,\mathbf{u}_{N_{e}}]^{T}$ is the global vector of all modal coefficients.
The stability of the fully discrete scheme depends on the interaction between spatial discretization and time integration. We employ arbitrary high order Deferred Correction (DeC) schemes [15, 22] of order $p+1$ for time marching. For a linear system like (10), the amplification factor of a $(p+1)$-th order DeC scheme can be expressed as
@@
where $\mathbf{U}^{n}=\mathbf{U}(t^{n})$, $\lambda$ denotes the eigenvalue of semi-discrete operator $\mathbf{M}^{-1}\mathbf{K}$, and $\mu=\lambda\Delta t$. In this study, spatial polynomial degrees up to $p=6$ are investigated, coupled with seventh-order DeC time integration. Consequently, the stability region in the complex $\mu$-plane is determined by the condition $|\sum_{k=0}^{p+1}\mu^{k}/k!|\leq 1$.
To perform the analysis is therefore needed to recast the ROD embedded boundary treatment in the general matrix form $K\mathbf{u}$. Although this is trivial for the SB polynomial correction as it was previously shown in [6] and recalled in section 4, this is not straightforward for the minimization-based ROD techniques which need careful manipulation to be written in such a form. In section 5, the ROD minimization problem is explicitly solved to show that this class of methods as well can be written as a special polynomial correction. This allows one to make a link between the two families of methods and, more interestingly allows one to perform a stability analysis.
@@ 3 Notation of embedded boundary methods
Figure 1: One-dimensional domain with an embedded boundary condition (red circle) imposed at a distance $d$ from $x_{L}$, which is defined as the computational boundary in this simplified configuration.
The work analyzes the stability of the ROD minimization-based treatment in the standard unfitted configuration where the physical boundary is separated from the computational mesh by a distance $d$ of the order of the mesh size $\Delta x$. This scenario is typical for immersed boundary simulations that employ a Cartesian background grid, as illustrated in figure 1. While in multi-dimensional settings the correspondence between a point on the true boundary and its projection onto the computational boundary may be defined in various ways (e.g., closest-point or level-set), the one-dimensional case admits a unique and straightforward mapping. This simplification permits a focused investigation of the stability properties introduced by the minimization-based boundary treatment.
@@
At the computational boundary, the embedded condition is enforced via a numerical flux. For the upwind flux considered here, the left-boundary flux is $\hat{u}(v_{h},u_{h}^{+})=v_{h}$, where $v_{h}$ represents the high order boundary reconstruction provided by the ROD treatment.